# Decoding Gray Matter: large-scale analysis of brain cell morphometry to inform microstructural modeling of diffusion MR signals

### v2

@@ 2.3 Shape descriptors
Many diffusion models of the diffusion MRI signal rely on modelling the cellular structure as a collection of randomly orientated cylinders in space [23, 51]. Whilst this was initially applied to modelling the signal from white matter where axons can be simplified to a collection of cylinders/sticks this model has been shown to apply to the dendrites of neurons [28, 29] and has been incorporated into current grey matter models of the diffusion signal.The cellular structure is modeled as a collection of cylinders, enabling the computation of fractional anisotropy (FA).
To calculate FA we followed the method outlined in [40]. This involves decomposing the cell into its primary projections and further segmenting these into cylinders of length $l=10\mu m$. From these segments the FA can be calculated by first computing the scatter matrix of the weighted line segments (weighted by their volume), from which the eigenvalues ($\tau_{i}$) can be found (Fig.3). The FA can be calculated from these eigenvalues using the following equation.
@@
It should be noted, due to the limited depth of field of some of the cellular reconstructions due to the method of acquisition, some cells appear flattened in dimension the perpendicular to the plane of the acquisition, which would result in an artificially high FA. To account for this, assuming that cells are axially symmetric, if the first eigenvalue was significantly lower than the second, $\tau_{1}<<\tau_{2}$, then the first eigenvalue was set to equal second, $\tau_{1}=\tau_{2}$, and all eigenvalues were normalized such that, $\sum_{i=1}^{3}\tau_{i}=1$.
The FA was calculated for all cells and the mean and standard deviation are reported.
@@
In addition to the FA of the line segments, their orientation dispersion (OD) was also computed. OD provides additional insight into the degree of anisotropy in cellular structures, complementing FA by describing the variability in the directions of the neuronal projections. The variability is typically characterised through a Watson distribution, which is a probability distribution of orientations around the primary axis on the unit sphere [52]. With the degree of clustering defined by the concentration parameter, $\kappa$. This concentration parameter can be used to calculate the OD through the following equation [23].
@@
Figure 4: A comparison between two cell types, mouse/rat pyramidal and granule cells. Showing exemplar cells over layed with decomposed line segments, the line segments centered at the origin, and orientation distribution about the z axis of the line segments and the analytical distribution given the calculated Watson concentration parameter
For cases where the data is not axially symmetric, such as cellular reconstructions with limited depth, the Watson distribution is insufficient as the orientation data is highly anisotropic. Instead, we applied the more general Bingham distribution, which accounts for orientation variability along multiple axes. The Bingham distribution models anisotropic orientation data and provides two concentration parameters ( $\kappa_{1}$ and $\kappa_{2}$ ), corresponding to the clustering along two orthogonal directions/planes.
To account for this variability in the cellular reconstructions a Bingham distribution was fitted to the weighted line segments using the MATLB tool box libDirectional [53] (Fig.4). From the resulting concentration parameters, if the first concentration parameter was significantly larger than the second $\kappa_{1}>>\kappa_{2}$, indicating the orientation data was highly planar, $\kappa_{2}$ was used as the Watson distribution parameter $\kappa$. Otherwise, the average of $\kappa_{1}$ and $\kappa_{2}$ was used as the Watson distribution parameter $\kappa$.
@@ 2.4 Topological descriptors
Another way of characterising cells is to look at their topology, which can provide a valuable means of characterising their complex branching structures. One such characterisation is the Topological Morphology Descriptor (TMD) [54]. The TMD returns the topological persistence barcode of the given structure. This barcode concisely encodes the branching structure of the cell, and has been shown to be highly effective in the generation of synthetic neurons [55]. Here, we apply the TMD to characterize the cellular projections and compare cellular topology across cell-types and species.
@@
Topological persistence analyses at what length scales a given topological feature, here the path length of connected branches, persists. The TMD is computed by tracking the initiation points and termination points, with respect to path length from the soma, of the connected branches at different length scales [56], preserving the longer components and filtering out the shorter ones. This is computed over the individual neurite projections, not the entire cell.
The terminal branches are evaluated with respect to their path length from the soma. Branches that share a parent node are compared and the branch with the shorter path length is filtered and its initiation length (path length of the parent node) and termination length are recorded. The longer branch is preserved and the process is repeated for all remaining branches, until there is only one remaining branch that corresponds to the path from the soma to the terminal point with the greatest path length. The returned barcode is a multi set of pairs of numbers, describing the initiation and termination of each constituent bar in the neurite, with respect to the path distance fro the soma. The topological barcode has been shown to retain detailed information of the structure it was computed for and can be used to categorise and identify cell types [55].
In this work the cells were first decomposed into their principle projections, by identifying the projections originating from the soma, and persistence barcodes from these projections were identified. The identified barcodes were used to create persistence images for each cell type (Fig.5). These images visualize the persistence of components based on their birth and death lengths. To enhance smoothness, Gaussian smoothing was applied, and the images were normalized so that the sum of their entries equaled one
Comparisons were made between persistence images by computing the global topological distance, D, between images,
$D=\frac{\sum_{i=1}^{n}|Im1_{i}-Im2_{i}|}{n}$
This topological distance was calculated for cell types with in species and also cell types between species.
@@ 3.1 Morphological features’ reference values
A summary of the structural features typical values can be seen in Tab.2. Some features show little variation between cell type and species ($<R_{branch}>_{s}$, $CV_{branch}$, $<\mu OD_{branch}>_{s}$, and $\tau_{branch}$). The remaining features displayed a wide range of values, suggesting higher inter-cellular and intra-species variability.
@@ 3.2 Comparing neuronal and glial cells
Given the growing interest in differentiating neurons and glial cells, the mean values of the structural features were computed across only neuronal and glial cells and reported in Tab.3. The findings show, compared to glial cells, neurons had larger soma and cell domain and longer branches; and reduced branching, branch curvedness, number of primary projections, and proportion of soma surface covered by projections; the remaining features displayed similar values.
@@ 3.3 Cellular shape reference values
A summary of the shape descriptors typical values for each cell type and species are reported in Tab.4. Cells with highly oriented and polarized projections, such as Purkinje cells, granule cells and pyramidal neurons have high FA and low orientation dispersion; while most of the glial cells projections are highly dispersed and with low FA.
Cell Type Species $N$ $R_{domain}(\mu m)$ $N_{proj}$ $BO$ $S/V_{domain}$ $R_{soma}(\mu m)$ $\eta_{soma}(\%)$ $S/V_{soma}(\mu m^{-1})$ $S/V_{branch}(\mu m^{-1})$ $L_{branch}(\mu m)$ $R_{branch}(\mu m)$ $CV_{branch}$ $\mu OD_{branch}$ $R_{c}(\mu m)$ $\tau_{branch}$
mouse/rat 169(123) $62\pm 70$ $9\pm 3$ $9\pm 4$ $2.5\pm 1.1$ $\geq 6.1\pm 6.3$ $\leq 4.7\pm 2.9$ $0.52\pm 0.38$ $4.3\pm 1.6$ $20\pm 13$ $\geq 2.1\pm 2.5$ 0.95 $0.26\pm 0.05$ $\geq 34\pm 31$ $0.89\pm 0.03$
Microglia monkey 61(60) $42\pm 8$ $8\pm 2$ $10\pm 4$ $5.2\pm 1.1$ $2.73\pm 0.50$ $\leq 1.4\pm 1.1$ $1.0\pm 0.25$ $16\pm 2$ $7.2\pm 5.7$ $geq0.29\pm 0.19$ 1.2 $0.30\pm 0.04$ $0.91\pm 0.02$ $0.91\pm 0.02$
human n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
mouse/rat 269(247) $40\pm 14$ $11\pm 7$ $19\pm 8$ $2.3\pm 0.65$ $\geq 2.1\pm 0.96$ $\leq 86\pm 26$ $0.88\pm 0.37$ $6.3\pm 2.2$ $4.9\pm 2.6$ $\geq 2.4\pm 1.0$ 0.55 $0.26\pm 0.05$ $\geq 3.5\pm 1.2$ $0.88\pm 0.03$
Astrocyte monkey n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
human n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
mouse/rat 80(80) $71\pm 19$ $15\pm 5$ $10\pm 3$ $0.67\pm 0.24$ $\geq 8.9\pm 1.7$ $\leq 61\pm 2.0$ $0.35\pm 0.06$ $13\pm 0.1$ $12.4\pm 3.4$ $\geq 0.3\pm 0.0$ 0.62 $0.16\pm 0.02$ $10\pm 1.1$ $0.81\pm 0.05$
Oligodendrocyte monkey n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
human n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
mouse/rat 351(330) $427\pm 350$ $5\pm 3$ $9\pm 4$ $3.3\pm 2.1$ $\geq 7.1\pm 3.6$ $\leq 18\pm 30$ $0.38\pm 0.18$ $12\pm 13$ $64\pm 30$ $\geq 1.0\pm 1.2$ 1.33 $0.26\pm 0.12$ $\geq 97\pm 173$ $0.86\pm 0.09$
Pyramidal monkey 867(858) $480\pm 248$ $6\pm 4$ $8\pm 3$ $2.7\pm 1.0$ $\geq 6.2\pm 3.7$ $\leq 24\pm 39$ $0.39\pm 0.17$ $5.7\pm 3.0$ $99\pm 75$ $\geq 1.1\pm 1.3$ 1 $0.23\pm 0.09$ $\geq 65\pm 62$ $0.89\pm 0.06$
human 1076(1070) $287\pm 95$ $6\pm 2$ $7\pm 2$ $1.9\pm 1.4$ $\geq 10\pm 3.6$ $\leq 7.2\pm 5.0$ $0.27\pm 0.09$ $5.8\pm 2.4$ $70\pm 21$ $\geq 1.5\pm 2.2$ 0.95 $0.29\pm 0.06$ $\geq 43\pm 21$ $0.90\pm 0.05$
mouse/rat 128(124) $358\pm 124$ $2\pm 1$ $5\pm 2$ $3.6\pm 1.0$ $\geq 4.1\pm 0.8$ $\leq 37\pm 19$ $0.68\pm 0.17$ $6.7\pm 2.6$ $109\pm 31$ $\geq 0.80\pm 0.050$ 0.51 $0.30\pm 0.05$ $\geq 16\pm 9$ $0.86\pm 0.04$
Granule monkey n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
human 77(77) $538\pm 110$ $5\pm 1$ $5\pm 1$ $2.8\pm 1.0$ $\geq 5.3\pm 2.7$ $\leq 77\pm 20$ $0.76\pm 0.38$ $3.6\pm 0.9$ $122\pm 27$ $\geq 1.0\pm 0.2$ 0.24 $0.20\pm 0.04$ $\geq 17\pm 8$ $0.84\pm 0.05$
mouse/rat 121(120) $158\pm 37$ $2\pm 1$ $10\pm 2$ $1.7\pm 0.49$ $\geq 7.6\pm 1.3$ $\leq 24\pm 18$ $0.36\pm 0.07$ $4.7\pm 2.2$ $8.6\pm 1.5$ $\geq 1.2\pm 0.7$ 0.43 $0.27\pm 0.05$ $\geq 21\pm 9$ $0.95\pm 0.01$
Purkinje monkey n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
human n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
mouse/rat 16(13) $449\pm 242$ $7\pm 2$ $10\pm 6$ $2.5\pm 0.65$ $\geq 6.3\pm 3.0$ $\leq 35\pm 37$ $0.43\pm 0.15$ $15\pm 11$ $57\pm 36$ $\geq 1.9\pm 3.8$ 1.2 $0.30\pm 0.12$ $\geq 11\pm 4$ $0.88\pm 0.06$
Basket monkey 20(18) $671\pm 302$ $9\pm 6$ $19\pm 7$ $2.5\pm 0.97$ $\geq 5.9\pm 2.0$ $\leq 3.7\pm 6.5$ $0.50\pm 0.18$ $12\pm 5$ $64\pm 23$ $\geq 0.7\pm 0.47$ 0.31 $0.35\pm 0.06$ $\geq 14\pm 6$ $0.79\pm 0.06$
human 25(21) $300\pm 344$ $5\pm 3$ $7\pm 6$ $3.2\pm 0.76$ $\geq 4.75\pm 0.65$ $\leq 24\pm 23$ $0.60\pm 0.12$ $7.7\pm 8.9$ $43\pm 39$ $\geq 2.6\pm 2.1$ 0.86 $0.24\pm 0.13$ $\geq 8.2\pm 5.0$ $0.90\pm 0.08$
mouse/rat 188(157) $284\pm 186$ $6\pm 2$ $11\pm 9$ $1.3\pm 0.29$ $\geq 7.6\pm 3.4$ $\leq 24\pm 24$ $0.25\pm 0.05$ $6.5\pm 2.6$ $57\pm 38$ $\geq 4.0\pm 3.1$ 0.98 $30\pm 0.07$ $\geq 47\pm 61$ $0.87\pm 0.04$
Gabaergic monkey n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
human 8(8) $350\pm 186$ $6\pm 3$ $11\pm 4$ $4.3\pm 0.64$ $\geq 4.3\pm 0.3$ $\leq 1.0\pm 0.3$ $0.65\pm 0.09$ $13.6\pm 2.7$ $52\pm 18$ $\geq 0.59\pm 0.36$ 1.7 $0.36\pm 0.07$ $\geq 12\pm 3$ $0.79\pm 0.05$
mouse/rat 36(35) $756\pm 675$ $6\pm 2$ $11\pm 6$ $1.7\pm 0.38$ $\geq 8.3\pm 2.6$ $\leq 6\pm 16$ $0.33\pm 0.05$ $5.4\pm 1.2$ $46\pm 17$ $\geq 1.6\pm 1.2$ 0.53 $0.36\pm 0.06$ $\geq 16\pm 7$ $0.85\pm 0.03$
Glutamatergic monkey n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
human 106(105) $616\pm 111$ $4\pm 2$ $6\pm 2$ $0.84\pm 1.7$ $\geq 29\pm 10$ $\leq 0.4\pm 3.3$ $0.10\pm 0.03$ $4.0\pm 0.1$ $145\pm 56$ $\geq 1.9\pm 1.5$ 5.8 $0.11\pm 0.03$ $\geq 475\pm 84$ $0.95\pm 0.02$
@@
Table 2: Summary of the morphological features computed for each species and cell type. N is the number of cellular structures investigated with complete information about the neurite structure; same information for soma in brackets. The reported values are mean $\pm$ s.d. over the corresponding sample. The $\geq$’ and $\leq$’ are used when the estimated value of the corresponding feature may be slightly (on average $<20\%$) under- or over-estimated, respectively, given the known limitations of the approach used. n.a. = not available.

### v4

@@ 2.3 Shape descriptors
Many diffusion MRI models represent cellular structure as a collection of randomly oriented cylinders [24, 58]. Initially applied to the signal from white matter where axons can be simplified to a collection of sticks this model has been shown to apply to the dendrites of neurons [32, 33] and has been incorporated into current grey matter diffusion MRI-based models.
To compute the fractional anisotropy (FA) for the cellular structure, the cellular structure was modeled as a collection of cylinders, following the method described in [45]. First, the cells were decomposed into primary projections and further segmented into cylinders of length $10\mu m$ (Fig.3). FA is calculated by first computing the scatter matrix of the weighted line segments orientations (weighted by volume), from which the eigenvalues ($\tau_{i}$) are derived. FA is then calculated as:
@@
In some cellular reconstructions, limitations in depth of field resulting from the imaging technique lead to anisotropic inaccuracies (see Supplementary Fig.2), with cells appearing compressed along the axis perpendicular to the acquisition plane (Z-axis). This artifact can result in overestimated fractional anisotropy in said reconstructions. For this reason, alongside the FA, we also report the estimated eigenvalues in ascending order $(\tau_{1}\leq\tau_{2}\leq\tau_{3})$ and an adjusted FA (Adj.FA) were we assumed $\tau_{1}=\tau_{2}$.
The mean and standard deviation are reported for all cell types.
@@
In addition to the FA of the line segments, their orientation dispersion (OD) was also computed. OD provides additional insight into the degree of anisotropy in cellular structures, complementing FA by describing the variability in the orientations of the neuronal projections. The variability is typically characterised through a Watson distribution, which is a probability distribution of orientations around the primary axis on the unit sphere [59]. With the degree of clustering defined by the concentration parameter, $\kappa$. This concentration parameter can be used to calculate the OD through the following equation [24].
@@
Figure 4: A comparison between two cell types, mouse/rat pyramidal and granule cells. Showing exemplar cells overlaid with decomposed line segments, the line segments centered at the origin, and orientation distribution about the z axis of the line segments and the analytical distribution given the calculated Watson concentration parameter
For cases where the data is not axially symmetric, such as cellular reconstructions with limited depth, the Watson distribution is insufficient. Instead, we applied the more general Bingham distribution, using MATLAB toolbox libDirectional [60] (Fig.4), which accounts for orientation variability along multiple axes. The Bingham distribution models anisotropic orientation data and provides two concentration parameters ( $\kappa_{1}$ and $\kappa_{2}$ ), corresponding to the clustering along two orthogonal directions.
From the resulting concentration parameters, if the first concentration parameter was significantly larger than the second $\kappa_{1}>>\kappa_{2}$, indicating the orientation data was highly planar, $\kappa_{2}$ was used as the Watson distribution parameter $\kappa$. Otherwise, the average of $\kappa_{1}$ and $\kappa_{2}$ was used as the Watson distribution parameter $\kappa$.
@@ 2.4 Topological descriptors
Another way to characterize cells is by analysing their topology, which captures the complexity of their branching structures. One approach to this is the Topological Morphology Descriptor (TMD) [61], which computes a topological persistence barcode for a given structure. This barcode provides a compact representation of the cell’s branching architecture and has proven effective in classifying neurons [61]. In this study, we use the TMD to quantify cellular projections and compare topological structures across different cell types and species.
Topological persistence quantifies how long specific structural features, such as branch paths, remain across different scales. The TMD captures this by measuring the path lengths of connected branches, tracking both their initiation and termination points relative to the soma [61]. This process preserves longer (more persistent) structural components while filtering out shorter ones.
@@
To compute a neuron’s topological descriptor, terminal points are evaluated by their path length from the soma. At each branching point, the shorter of the two sibling branches is removed, and its initiation and termination distances from the soma are recorded in the topological barcode. The longer branch is retained, and this process is repeated until all terminal points have been assessed. The resulting barcode is a multiset of value pairs, each representing the initiation and termination lengths of a branch segment relative to the soma. The longer branch is retained, and this process is repeated until all terminal points have been assessed. The resulting barcode is a multiset of value pairs, each representing the initiation and termination points with respect to the soma.
These barcodes retain rich structural information and can be used to categorize and identify cell types [62]. In our work, each cell was first decomposed into its principal projections by identifying dendritic projections emanating from the soma. Persistence barcodes were then computed for each projection and used to generate persistence images (Fig. 5), which visualize the distribution of persistent features by their initiation and termination lengths. To improve visual smoothness, kernel density estimation was applied, and each image was normalized so that the sum of its pixel values equaled one.
Comparisons were made between persistence images by computing the global topological distance [61], D, between images,
$D_{celltype1,celltype2}=\sum_{i=1}^{n}|celltype1_{i}-celltype2_{i}|$
where $celltype1$ and $celltype2$ are the two persistence images being compared. With D ranging from zero, the persistent images are identical and over lab completely, to two, the persistent images share no overlab or similarity. This topological distance was calculated for cell types within species and also cell types between species.
@@ 3.1 Morphological features’ reference values
A summary of the typical values of structure features is given in Tab.2. Some features show little variation between cell type and species ($R_{branch}>_{s}$, $CV_{branch}$, $\mu OD_{branch}>_{s}$, and $\tau_{branch}$). The remaining features displayed a wide range of values, suggesting higher inter-cellular and intra-species variability.
@@ 3.2 Comparing neuronal and glial cells
Given the growing interest in differentiating neurons and glial cells, the mean values of the structural features were computed across only neuronal and glial cells and reported in Tab.3. The findings show, compared to glial cells, neurons have larger soma and cell domain and longer branches; and reduced branching, branch curvedness, number of primary projections, and proportion of soma surface covered by projections; the remaining features displayed similar values.
3.3 Correlations between Morphological features
The correlations between structural features are illustrated in Supplementary Fig.4. Some consistent and expected patterns are observed, particularly involving surface-to-volume ratio measures: the soma surface-to-volume ratio is obviously negatively correlated with soma radius, and the branch surface-to-volume ratio is obviously negatively correlated with branch radius. Furthermore, the surface-to-volume ratio of both the soma and branches are obviously positively correlated with the surface-to-volume ratio of the domain. Additional, micro-orientation dispersion shows a positive correlation with branch tortuosity, $\tau$, across majority of cell types. No further consistent correlations between features were observed across cell types.
@@ 3.4 Cellular shape reference values
A summary of the shape descriptors typical values for each cell type and species is reported in Tab.4. Cells with highly oriented and polarized projections, such as Purkinje cells, granule cells and pyramidal neurons have high FA and adjFA and low orientation dispersion; while most of the glial cells projections are highly dispersed and with low FA and adjFA.
NN $R_{domain}$ $N_{proj}$ BOBO $S/V_{domain}(\mu m^{-1})$ $S/V_{soma}(\mu m^{-1})$ $R_{soma}(\mu m)$ $R_{MRsoma}(\mu m)$ $\eta_{soma}(\%)$ $S/V_{branch}(\mu m^{-1})$ $L_{branch}(\mu m)$ $R_{branch}(\mu m)$ $CV_{branch}$ $R_{MRbranch}(\mu m)$ $\mu OD_{branch}$ $R_{c}(\mu m)$ $\tau_{branch}$
celltype animal $Q1$ median $Q3$ $Q1$ median $Q3$ $Q1$ median $Q3$ $Q1$ median $Q3$ $Q1$ median $Q3$ $Q1$ median $Q3$ $Q1$ median $Q3$ $Q1$ median $Q3$ $Q1$ median $Q3$ $Q1$ median $Q3$ $Q1$ median $Q3$ $Q1$ median $Q3$ $Q1$ median $Q3$
MouseRat 4100 (4072) 33 41 49 3 5 6 5.6 7.8 9.7 2.6 3.3 3.6 0.10 1.3 2.0 1.6 2.5 3.1 3.5 4 48 99 2.3 3.4 6.2 6.4 7.4 9.8 0.67 1.0 1.6 0.58 1.0 0.19 0.2 0.25 3.4 3.9 7.5 0.77 0.81 0.82
Microglia Monkey 60 (60) 37 42 47 7 8 9 8.4 9.2 10 2.9 3.9 4.9 1 1.2 1.3 2.3 2.6 3.0 3 0.7 1.1 1.6 15 17 18 8.9 9.4 9.9 0.18 0.2 0.36 0.63 0.2 0.25 0.27 0.29 13 16 18 0.73 0.74 0.75
Human n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a
MouseRat 310 (307) 32 39 45 5 7 12 6.9 8 14 0.37 4.1 5.5 0.8 0.9 1.2 2.5 3.5 4.0 4.3 3 27 70 4.2 6.7 8.8 6.5 7.8 9.3 0.46 0.78 1.6 0.93 0.54 0.21 0.27 0.3 5.1 9.6 14 0.66 0.69 0.72
Astrocyte Monkey n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a
Human n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a
MouseRat 58 (58) 85 100 130 5 7 9 7 7.9 9 2.8 3 3.1 0.61 0.74 0.90 3.4 4.1 5.0 5 4 7 13 3.9 4.8 5.5 22 28 32 0.57 0.82 1.3 1.20 0.64 0.15 0.16 0.21 21 27 34 0.77 0.81 0.83
Oligodendrocyte Monkey n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a
Human n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a
MouseRat 3005 (3000) 170 290 490 4 6 7 6.4 7.9 10 0.34 0.34 0.86 0.35 0.45 0.7 4.4 6.7 8.8 11 1 3 10 3.2 4.6 7.4 37 52 72 0.48 0.84 1.3 0.63 0.99 0.17 0.23 0.29 14 26 41 0.76 0.8 0.84
Pyramidal Monkey 510 (508) 250 370 630 6 7 9 7.2 8.2 9.6 2.5 2.9 3.3 0.37 0.43 0.5 6.1 7.2 8.3 8.5 0.2 3.2 8.1 4.5 6.3 9.9 53 62 72 0.44 0.69 1.2 0.74 0.79 0.18 0.22 0.29 22 39 54 0.75 0.8 0.83
Human 2403 (2400) 230 270 310 5 6 7 6.2 6.7 7.4 3 3.3 3.5 0.32 0.37 0.42 7.2 8.3 9.4 9.7 4.6 6.9 9.6 4.1 5.2 6.6 64 74 86 0.61 0.69 1.2 0.58 0.88 0.24 0.27 0.3 49 63 79 0.72 0.75 0.78
MouseRat 930 (928) 100 160 230 1 1 1 3 3.8 4.5 1.6 2 2.7 0.64 0.73 0.85 3.6 4.2 4.8 5.1 1.0 1.8 3.8 3.2 4.1 5.7 27 45 78 0.63 0.93 1.3 0.45 0.78 0.12 0.16 0.21 23 35 56 0.8 0.84 0.87
Granule Monkey n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a
Human 82 (82) 460 530 580 2 3 4 4 5 6 2.1 2.6 3.3 0.41 0.77 1.0 3.1 4.0 7.4 7.9 4 8 15 3.6 4.2 4.5 110 130 150 0.53 0.64 0.75 0.49 0.76 0.18 0.20 0.24 15 18 23 0.83 0.85 0.87
MouseRat 140 (140) 130 160 190 1 2 2 9.3 10 12 2.2 2.5 2.7 0.36 0.41 0.46 6.6 7.4 8.3 8.3 2 26 33 3.4 4.3 5.3 8.9 9.9 11 0.76 0.99 1.4 0.22 0.7 0.25 0.28 0.31 17 25 32 0.73 0.78 0.81
Purkinje Monkey n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a
Human n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a
MouseRat 441 (440) 190 240 300 4 5 7 5 6 7 1.5 2.0 2.7 0.39 0.51 0.66 4.6 5.9 7.8 9.1 1 3 38 2.7 4.0 5.6 50 64 81 0.49 0.61 0.94 0.77 0.93 0.22 0.27 0.33 11 23 45 0.73 0.77 0.81
Basket Monkey n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a
Human 11 (11) 220 270 380 4 5 6 4.5 6 6.9 3 3 3.2 0.56 0.58 0.67 4.6 5.2 5.4 5.6 7 8 45 3.4 4 4.6 51 62 110 0.92 0.95 1.5 0.49 0.67 0.21 0.23 0.27 14 16 22 0.77 0.79 0.82
MouseRat 180 (179) 180 250 440 4 6 8 6.1 7.3 8.5 1.8 2.4 3 0.38 0.47 0.53 5.7 6.5 8.1 8.3 1 1.5 4.8 4.3 6.1 13 35 46 65 0.34 0.67 1.2 0.80 0.67 0.22 0.29 0.33 9.8 15 50 0.73 0.78 0.81
Gabaergic Monkey 4 (4) 120 250 370 6 7.5 10 5.5 6.5 7.1 2.6 3.2 3.6 0.55 0.62 0.68 4.7 5.1 5.8 5.5 0.27 0.27 0.27 4.9 5.5 6.7 35 67 120 0.72 0.76 0.93 0.51 0.44 0.3 0.34 0.37 2.1 2.3 2.5 0.7 0.72 0.76
Human 6 (6) 140 190 220 2 3 5 5.2 5.4 5.6 4 4.1 4.5 0.67 0.69 0.72 4.2 4.5 4.6 4.5 0.38 0.75 0.9 7.5 7.8 8 63 65 83 0.48 0.5 0.67 0.57 0.38 0.28 0.30 0.33 13 17 27 0.73 0.74 0.74
MouseRat 106 (106) 150 250 360 3 5 7 5 7 8 1.2 2.7 2.3 0.34 0.42 0.50 6.1 7.3 9.0 9.7 1 1.8 7.2 3.9 5.1 20 42 54 72 0.21 0.63 0.80 0.73 0.98 0.21 0.27 0.34 20 27 56 0.71 0.81 0.84
Glutamatergic Monkey n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a
Human n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a n.a
@@
Table 2: Summary of the morphological features computed for each species and cell type. N is the number of cellular structures investigated with complete information about the neurite structure; same information for soma in brackets. The reported values are mean $\pm$ s.d. over the corresponding sample. The $\geq$’ and $\leq$’ are used when the estimated value of the corresponding feature may be slightly (on average $20\%$) under- or over-estimated, respectively, given the known limitations of the approach used. n.a. = not available.